# The Artificial Benchmark for Community Detection with Outliers and Overlapping Communities (ABCD+o2)

### v1

@@ 2.5 The ABCD+o2 Construction
Before detailing the pairing process, let us explain how we deal with the correlation parameter $\rho$. Recall that our goal is to pair the $\hat{n}$ non-outlier degrees with the $\hat{n}$ elements such that the Pearson correlation between the degree and number of communities is approximately $\rho$. To this end, we perform multiple pairing attempts, with each attempt aiming to achieve a closer approximate the desired correlation. In each attempt, we fix a parameter $\alpha\in\mathbb{R}$ with negative values of $\alpha$ producing negative correlations and positive values producing positive correlations. The parameter $\alpha$ is tuned to achieve a correlation close to the desired value $\rho$. This is achieved with a binary search, starting with $\alpha_{\min}=-60$ and $\alpha_{\max}=60$ and stopping once the desired precision is achieved or after the obtained correlation level stops improving. In some cases it is impossible to reach the correlation level required by the user, in which case the closest possible correlation is produced.
In each pairing attempt with parameter $\alpha\in\mathbb{R}$, we iteratively assign degrees of non-outliers to elements as follows. Recall that the degree sequence $\hat{\textbf{d}}_{\hat{n}}$ is sorted with $\hat{d}_{1}$ being the maximum degree. Starting with $i=1$, let $R_{i}$ be the collection of unassigned elements at step $i$. At step $i$, choose element $v$ with probability proportional to $\eta_{v}^{\alpha}$ from the set of elements in $R_{i}$ satisfying
$\hat{d}_{i}\leq\frac{\eta_{v}}{1-\xi\phi}\cdot\min\Big{\{}|C_{k}|-1:j\in C_{k}\Big{\}}\,,$ (3)
where recall that $\eta_{v}$ is the number of communities element $v$ belongs to. The chosen element $v$ gets paired with the $i$th degree in the subsequence $\hat{\mathbf{d}}_{\hat{n}}$, namely, the degree $\hat{d}_{i}$. If the set defined by the condition (3) is empty, then we sample from the elements in $R_{i}$ for which the value of $({\eta_{v}}/({1-\xi\phi}))\cdot\min\{|C_{k}|-1:j\in C_{k}\}$ maximal.
Once the pairing is complete, we relabel as follows. The nodes are labelled as $[n]$, with node $i$ having degree $d_{i}$. The communities $(C_{j},j\in L)$ now contain non-outlier nodes corresponding to degree/element pairing. If $d_{i}$ was assigned a non-outlier, its corresponding node $i$ belongs to some number $\eta_{i}$ of communities, one of which is its primary community. If $d_{i}$ was assigned to an outlier, its corresponding node $i$ is assigned to an auxiliary “community” labelled $C_{0}$; the set $C_{0}$ is merely a label for the outliers and such an outlier $i$ has $\eta_{i}=0$.
@@
For each non-outlier $i\in[n]$ we split the $d_{i}$ half-edges of $i$ into community half-edges and background half-edges. To this end, define $Y_{i}:=\left\lfloor(1-\xi)d_{i}\right\rceil$ and $Z_{i}:=d_{i}-Y_{i}$ (note that $Y_{i}$ and $Z_{i}$ are random variables with ${\mathbb{E}}\left[Y_{i}\right]=(1-\xi)d_{i}$ and ${\mathbb{E}}\left[Z_{i}\right]=\xi d_{i}$) and, for all non-outliers $i\in[n]$, split the $d_{i}$ half-edges of $i$ into $Y_{i}$ community half-edges and $Z_{i}$ background half-edges. Community half-edges are further split into the $\eta_{i}$ communities the non-outlier node $i$ belongs to, as evenly as possible. Specifically, for the communities containing node $i$, $Y_{i}-\eta_{i}\lfloor Y_{i}/\eta_{i}\rfloor$ communities (chosen randomly) each receive $\lfloor Y_{i}/\eta_{i}\rfloor+1$ half-edges and the remaining communities each receive $\lfloor Y_{i}/\eta_{i}\rfloor$ half-edges. On the other hand, if $i\in[n]$ is an outlier then we set $Z_{i}=d_{i}$.
Once the assignment of degrees is complete, for each $j\in[L]$, we independently construct the community graph $G_{n,j}$ as per the configuration model on node set $C_{j}$ and the corresponding degree sequence. In the event that the sum of degrees in a community is odd, we pick a maximum degree node $i$ in said community and decrease its community degree by one while increasing its background graph degree by one. Finally, construct the background graph $G_{n,0}$ as per the configuration model on node set $[n]$ and degree sequence $(Z_{i},i\in[n])$. Let $G_{n}=\bigcup_{0\leq j\leq n}G_{n,j}$ be the union of all graphs generated in this phase.
@@
As a result, after ignoring edges in the global recycle list, all community graphs are simple and the background graph is simple. However, as is the case in the original ABCD model, an edge in the background graph can form a multi-edge with an edge in a community graph. Another problem that might occur, specific to ABCD+o2 model, is that an edge from one community can form a multi-edge with an edge from a different but overlapping community. All of these problematic edges are added to the global recycle list. We merge all community graphs with the background graph. Finally, the global recycle list is transformed into a list of half-edges and new edges are created from it. We follow the same procedure as for the community graphs. However, we do not “give up” recycling and follow the process until all required edges are created. As the background graph is sparse, this final rewiring is very fast in practice.
@@ 3 Properties of the ABCD+o2 Model
In this section, we present experiments highlighting the properties of the ABCD+o2 model. We begin by investigating the communities generated by the reference layer as described in Section 3.1. We then continue with additional features of the model in Section 3.2. We are mainly interested in comparing the model’s properties with properties of real-world networks. Thus, for the model parameters, we use values derived from several networks that have known overlapping communities. The networks were presented in [36], and are available on the Stanford Network Analysis Project (SNAP) website [^4] https://snap.stanford.edu/data/index.html#communities.
@@
For each network, all parameters other than $\delta$ and $s$ are measured directly from the data. We set $\delta=5$ to avoid cases where nodes have no neighbours in their own community, and insist on $s\geq 10$. The power-law exponents $\gamma$ and $\beta$ are fit using maximum-likelihood estimation described in [11] and implemented in the Python package powerlaw [^5] https://github.com/jeffalstott/powerlaw[1]. In this fitting process, we set the minimum degree as $\delta=5$ and the minimum community size as $s=10$ for consistency. The level of noise $\xi$ is computed directly from the graphs as the fraction $|E_{b}|/|E|$ where $E_{b}$ is the set of edges where the incident nodes share no communities. The parameters for each network are shown in Table 2.
@@
$n$ Number of nodes $317{,}080$ $334{,}863$ $52{,}675$
$s_{0}$ Number of outliers $56{,}082$ $17{,}669$ 00
@@
$\delta$ Min degree at least $\delta$ $5$ $5$ $5$
$\Delta$ Max degree at most $\Delta$ $343$ $549$ $1{,}928$
@@
$S$ Max community size at most $S$ $7{,}556$ $53{,}551$ $3{,}001$
@@
Additionally, some of the experiments are affected by the dimension $d$ of the hidden reference layer. In such cases, we compare dimensions $2$, $8$, and $64$.
@@ 3.1 Effects of the Reference Layer
The new and novel feature of the ABCD+o2 model is the hidden reference layer used to generate overlapping communities. Here, we analyze the effects of this reference layer on the community overlap sizes. For comparison, we consider the CKB model [10] following a description from [30]. The CKB model is an application of the random bipartite community affiliation graph proposed for overlapping LFR [25] that uses a power-law distribution for the number of communities per node. Let us briefly summarize this graph construction process. The community affiliation graph is a bipartite graph where one part consists of nodes and the other consists of communities. As input, the CKB model requires the number of nodes $n$, a truncated power-law distribution $\mathcal{P}\left(\Omega,x_{\min},x_{\max}\right)$ for the number of communities per node, and another truncated power-law distribution $\mathcal{P}\left(\beta,s,S\right)$ for the community sizes. The number of communities is set to
@@
In the first experiment (see Figure 2), we compare the distributions of community sizes. Of course, since both ABCD+o2 and CKB sample community sizes directly from the specified distribution, it is unsurprising that the empirical distributions are near-perfect fits. Thus, this first experiment merely acts as a sanity check ensuring that the ABCD+o2 and CKB models generate community sizes correctly. Moreover, this experiment verifies that the distributions coming from the real networks are indeed power-law.
Figure 2: Distribution of community sizes for the three real-world networks and their synthetic counterparts.
@@
Figure 3: Distribution of the number of communities per node.
@@
Figure 4: Distribution of the size of overlaps.
@@
YouTube $0.37$ $0.37$ $0.37$ $0.38$
Table 3: Measured $\rho$ in both the real networks and the corresponding ABCD+o2 graphs.
@@
One consequence of forcing nodes of large degree into community intersections is that those intersections will be denser than the individual communities. We show this phenomenon in Figure 5 by comparing the community and intersection densities for various $\rho$. Note that, given communities $C_{j}$ and $C_{k}$ with $s_{j}\gg s_{k}$, the density of $C_{k}$ is nearly always larger than that of $C_{j}$ since both community graphs are sparse (at most a $1-\xi$ fraction of each node’s degree is delegated to one of its communities). Figures for the other graphs and dimensions can be found in Appendix A. For computational reasons, we only consider overlaps of size at least 25, and to ensure we are not simply measuring the density of the smaller community, the size of the overlap must be at most half the size of the smaller community.
Figure 5: Community overlap densities compared to individual community densities for various values of $\rho$. The line corresponds to the median, and the shaded region is the 25th to 75th quantile. We only consider pairs of communities where the overlap is at least 50 nodes and the size of the overlap is less than half the size of the smaller community.
Having community intersections that are denser than the respective communities is a desirable property and one of our main motivators for introducing the parameter $\rho$ to the ABCD+o2 model. In [37] it was found that, in real networks, community intersections can be significantly denser than the individual communities. In Figure 6 we confirm that the DBLP, Amazon, and YouTube networks all exhibit this phenomenon. Moreover, we compare these densities with those coming from the corresponding ABCD+o2 graphs. Aside from confirming that the model mimics the behaviour of real networks, we also find that the dimension of the reference layer affects the density of both the communities and the intersections. Moreover, we see that a low-dimensional reference layer yields a better approximation of the real data in the case of Amazon, whereas a high-dimensional reference layer yields a better approximation in the cases of DBLP and YouTube. Note that this high-dimensional vs. low-dimensional discrepancy is consistent with the experiment presented in Figure 3.
Figure 6: Density of the Overlap compared to the density of the communities. We restrict to overlaps of size at least $10$ due to the discrete and unpredictable nature of small communities.
@@
Finally, we analyze how strongly nodes are connected to their communities. There are many measures for community association strength, some of which we recently analyzed in [4]. For the sake of simplicity, the measure we will use here is the internal edge fraction $\mathrm{IEF}(v,C)$, defined as
@@
Figure 7: Five largest internal edge fractions, sorted in descending order, grouped by the number of communities.
@@ 4 Benchmarking Community Detection Algorithms
Leiden [33]: a greedy algorithm that attempts to optimize the modularity function. Note that this algorithm returns a partition, and we use it merely as a baseline to compare with algorithms that attempt to find overlapping communities.
Edge Clustering [24]: an edge-partitioning algorithm that translates to overlapping clusters of nodes. Pairs of edges are measured based on similarity of neighbourhoods, and these similarity measures dictate the order in which edge-communities merge. As edge-communities merge, the modularity is tracked on the line-graph, and the maximum modularity attained yields the edge-communities, which in turn yields overlapping node-communities.
Ego-Split [13]: a method which finds overlapping clusters in a graph $G$ by applying a partitioning algorithm such as Leiden to an auxiliary graph $G^{\prime}$ and mapping the resulting partition onto $G$. The auxiliary graph $G^{\prime}$ is constructed from $G$ by creating multiple copies, or “egos”, of each node based on its neighbourhood.
Ego-Split+IEF [4]: the same algorithm as Ego-Split, but with a post-processing step that re-assigns nodes to communities based on the IEF measure.
@@
In the first experiment, we compare the accuracy of each algorithm with respect to the ground-truth communities. The measure we use to determine accuracy is the overlapping Normalized Mutual Information (oNMI) measure [27]. For this experiment we fix the number of nodes $n=5,000$, the minimum degree $\delta=5$, the maximum degree $\Delta=100$, the minimum community size $s=50$, and the maximum community size $S=500$. The parameters of ABCD+o2 with the most influence on the quality of detection algorithms are $\xi$ (the level of noise) and $\eta$ (the average number of communities a non-outlier is part of), and we consider every combination of $\xi\in\{0.1,0.2,\dots,0.6\}$ and $\eta\in\{1,1.25,\dots,2.5\}$. We fix the remaining parameters according to the YouTube graph, namely, $\gamma=1.87$, $\beta=2.13$, and $\rho=0.37$.
@@
Figure 10: The performance of four clustering algorithms on ABCD+o2 graphs with 1,000 to 15,000 nodes.

### v2

@@ 2.5 The ABCD+o2 Construction
Before detailing the pairing process, let us explain how we deal with the correlation parameter $\rho$. Recall that our goal is to pair the $\hat{n}$ non-outlier degrees with the $\hat{n}$ elements such that the Pearson correlation between the degree and number of communities is approximately $\rho$. To this end, we perform multiple pairing attempts, with each attempt aiming to achieve a closer approximate the desired correlation. In each attempt, we fix a parameter $\alpha\in\mathbb{R}$ with negative values of $\alpha$ producing negative correlations and positive values producing positive correlations. Once the pairing is finished, we compute the obtained correlation $\rho_{\alpha}$. The parameter $\alpha$ is tuned to achieve correlation $\rho_{\alpha}$ close to the desired value $\rho$. This is achieved with a binary search, starting with $\alpha_{\min}=-60$ and $\alpha_{\max}=60$ and stopping once the desired precision is achieved or after the obtained correlation level stops improving. In some cases it is impossible to reach the correlation level required by the user, in which case the closest possible correlation is produced.
[…]
$\hat{d}_{i}\leq\frac{\eta_{v}}{1-\xi\phi}\cdot\min\Big\{|C_{k}|-1:v\in C_{k}\Big\}\,$, (3)
where recall that $\eta_{v}$ is the number of communities element $v$ belongs to. As a result, as desired, parameter $\alpha$ can be tuned to get negative correlation (negative $\alpha$) as well as positive one (positive $\alpha$) but the relationship between $\alpha$ and the correlation parameter $\rho$ is unclear. The chosen element $v$ gets paired with $\hat{d}_{i}$. If the set defined by the condition (3) is empty, then we sample $v$ from the elements in $U_{i}$ for which the value of $({\eta_{v}}/({1-\xi\phi}))\cdot\min\{|C_{k}|-1:v\in C_{k}\}$ is maximal.
Once the pairing is complete, we relabel as follows. The nodes are labelled as $[n]$, with node $i$ having degree $d_{i}$ (note that the inequality $d_{1}\geq\dots\geq d_{n}$ still holds). The communities $(C_{j},j\in L)$ now contain non-outlier nodes corresponding to degree/element pairing. If $d_{i}$ was assigned a non-outlier, its corresponding node $i$ belongs to some number $\eta_{i}$ of communities, one of which is its primary community. If $d_{i}$ was assigned to an outlier, its corresponding node $i$ is assigned to an auxiliary “community” labelled $C_{0}$; the set $C_{0}$ is merely a label for the outliers and such an outlier $i$ has $\eta_{i}=0$.
@@
For each non-outlier $i\in[n]$ we split the $d_{i}$ half-edges of $i$ into community half-edges and background half-edges. To this end, define $Y_{i}:=\left\lfloor(1-\xi)d_{i}\right\rceil$ and $Z_{i}:=d_{i}-Y_{i}$ (note that $Y_{i}$ and $Z_{i}$ are random variables with ${\mathbb{E}}\left[Y_{i}\right]=(1-\xi)d_{i}$ and ${\mathbb{E}}\left[Z_{i}\right]=\xi d_{i}$) and, for all non-outliers $i\in[n]$, split the $d_{i}$ half-edges of $i$ into $Y_{i}$ community half-edges and $Z_{i}$ background half-edges. Community half-edges are further split into the $\eta_{i}$ communities the non-outlier node $i$ belongs to, as evenly as possible. Specifically, for the communities containing node $i$, $Y_{i}-\eta_{i}\lfloor Y_{i}/\eta_{i}\rfloor$ communities (chosen randomly) each receive $\lfloor Y_{i}/\eta_{i}\rfloor+1$ half-edges and the remaining communities each receive $\lfloor Y_{i}/\eta_{i}\rfloor$ half-edges. Note that condition (3) makes sure that after splitting the community half-edges into communities, as evenly as possible, there is enough room in each community. On the other hand, if $i\in[n]$ is an outlier then we set $Z_{i}=d_{i}$.
Once the assignment of degrees is complete, for each $j\in[L]$, we independently construct the community graph $G_{n,j}$ as follows. For each node $i\in C_{j}$, there is a set of half-edges attached to $i$ delegated to $C_{j}$. With these delegated half-edges, we form $G_{n,j}$ as per the configuration model (see Subsection 2.2 for the definition of the model and relevant references). In the event that the sum of degrees in a community is odd, we pick a maximum degree node $i$ in said community and decrease its community degree by one while increasing its background graph degree by one. Finally, construct the background graph $G_{n,0}$ as per the configuration model on node set $[n]$ and degree sequence $(Z_{i},i\in[n])$. Let $G_{n}=\bigcup_{0\leq j\leq n}G_{n,j}$ be the union of all graphs generated in this phase.
@@
As a result, after ignoring edges in the global recycle list, all community graphs are simple and the background graph is simple. However, as is the case in the original ABCD model, an edge in the background graph can form a multi-edge with an edge in a community graph. Another problem that might occur, specific to ABCD+o2 model, is that an edge from one community can form a multi-edge with an edge from a different but overlapping community. All of these problematic edges are added to the global recycle list. We merge all community graphs with the background graph. Finally, the global recycle list is transformed into a list of half-edges and new edges are created from it. We follow the same procedure as for the community graphs. As the background graph is sparse, this final rewiring is very fast in practice. However, in rare cases it is possible that the process does not terminate. We try rewiring 100 times the size of the global recycle list. If, after this many iterations, we cannot find any new admissible edge (thus reducing the size of global recycle list) we throw an error and indicate to the user that the algorithm failed to terminate with success for the given input parameters (indicating that the likely cause is that the required input degree sequence was not graphic and it should be reconsidered by the user).
@@ 3 Properties of the ABCD+o2 Model
In this section, we present experiments highlighting the properties of the ABCD+o2 model. We begin by investigating the communities generated by the reference layer as described in Section 3.1. We then continue with additional features of the model in Section 3.2. We are mainly interested in comparing the model’s properties with properties of real-world networks. Thus, for the model parameters, we use values derived from several networks that have known overlapping communities. The networks were presented in [yang2012defining], and are available on the Stanford Network Analysis Project (SNAP) website [^4] https://snap.stanford.edu/data/index.html#communities.
@@
For each network, all parameters other than $\delta$ and $s$ are measured directly from the data. We set $\delta=5$ to avoid cases where nodes have no neighbours in their own community, and insist on $s\geq 10$. The power-law exponents $\gamma$ and $\beta$ are fit using maximum-likelihood estimation described in [clauset2009powerlaw] and implemented in the Python package powerlaw [^5] https://github.com/jeffalstott/powerlaw[alstott2014powerlaw]. In this fitting process, we set the minimum degree as $\delta=5$ and the minimum community size as $s=10$ for consistency. The level of noise $\xi$ is computed directly from the graphs as the fraction $|E_{b}|/|E|$ where $E_{b}$ is the set of edges where the incident nodes share no communities. The parameters for each network are shown in Table 2. Additionally, most of the experiments are affected by the dimension $d$ of the hidden reference layer. In such cases, we compare dimensions $2$, 88, and $64$.
@@
$n$ Number of nodes 317,080317{,}080 334,863334{,}863 52,67552{,}675
$s_{0}$ Number of outliers 56,08256{,}082 17,66917{,}669 0
@@
$\delta$ Min degree at least $\delta$ 55 55 55
$\Delta$ Max degree at most $\Delta$ $343$ $549$ 1,9281{,}928
@@
$S$ Max community size at most $S$ 7,5567{,}556 53,55153{,}551 3,0013{,}001
@@ 3.1 Effects of the Reference Layer
The new and novel feature of the ABCD+o2 model is the hidden reference layer used to generate overlapping communities. Here, we analyze the effects of this reference layer on the community overlap sizes. For comparison, we consider the CKB model [chykhradze2014ckb] following a description from [sengupta2017ckbdescription]. The CKB model is an application of the random bipartite community affiliation graph proposed for overlapping LFR [lancichinetti2009benchmarks] that uses a power-law distribution for the number of communities per node. Let us briefly summarize this graph construction process. The community affiliation graph is a bipartite graph where one part consists of nodes and the other consists of communities. As input, the CKB model requires the number of nodes $n$, a truncated power-law distribution $\mathcal{P}\left(\Omega,x_{\min},x_{\max}\right)$ for the number of communities per node, and another truncated power-law distribution $\mathcal{P}\left(\beta,s,S\right)$ for the community sizes. The number of communities is set to
@@
In the first experiment (see Figure 2), we compare the distributions of community sizes. Of course, since both ABCD+o2 and CKB sample community sizes directly from the specified distribution, it is unsurprising that the empirical distributions are near-perfect fits when accounting for the imposed minimum community size of $10$. Thus, this first experiment merely acts as a sanity check ensuring that the ABCD+o2 and CKB models generate community sizes correctly. Moreover, this experiment verifies that the distributions coming from the real networks are indeed power-law.
Figure 2: Distribution of community sizes for the three real-world networks and their synthetic counterparts. The y-axis shows the empirical complementary cumulative distribution function, computed as $y=1-|\{C_{i}:|C_{i}|<x\}|\;/\;|\{C_{i}\}|$.
@@
Figure 3: Distribution of the number of communities per node. The y-axis shows the empirical complementary cumulative distribution function, computed as $y=1-|\{v_{i}:|\{C_{j}:v_{i}\in C_{j}\}<x\}|\;/\;|V|$.
@@
Figure 4: Distribution of the size of overlaps. The y-axis shows the empirical complementary cumulative distribution function of non-empty overlaps. For example, the center column of figures corresponds to three overlapping communities and is computed as $y=1-|\{\{C_{i},C_{j},C_{k}\}:|C_{i}\cap C_{j}\cap C_{k}|<x\}|\;/\;|\{\{C_{i},C_{j},C_{k}\}:|C_{i}\cap C_{j}\cap C_{k}|\neq\emptyset|$.
@@
YouTube $0.37$ $0.37$ $0.36$ $0.38$
Table 3: Measured $\rho$ in both the real networks and the corresponding ABCD+o2 graphs. We generated 10 ABCD+o2 graphs for each dimension, and found that across all graphs and dimensions, the standard error, to two decimal places, was 0.
@@
One consequence of forcing nodes of large degree into community intersections is that those intersections will be denser than the individual communities. We show this phenomenon in Figure 5 by comparing the community and intersection densities for various $\rho$ while fixing the other parameters. Note that, given communities $C_{j}$ and $C_{k}$ with $s_{j}\gg s_{k}$, the density of $C_{k}$ is almost always larger than that of $C_{j}$ since both community graphs are sparse (at most a $1-\xi$ fraction of each node’s degree is delegated to one of its communities). Figures for the other graphs and dimensions can be found in Appendix A.
Figure 5: Community overlap densities compared to individual community densities for various values of $\rho$. For 9 ABCD+o2 graphs with $\rho$ evenly spaced in [0.5,0.5][-0.5,0.5], we compute the empirical value $\hat{\rho}$ and the distribution of density for overlapping communities. We plot $\hat{\rho}$ against the distribution of the density for the intersection and each community individually. The line corresponds to the median, and the shaded region to the 25th to 75th quantile. For computational reasons, we only consider overlaps of size at least 25, and to ensure we are not simply measuring the density of the smaller community, the size of the overlap must be at most half the size of the smaller community.
Having community intersections that are denser than the respective communities is a desirable property and one of our main motivators for introducing the parameter $\rho$ to the ABCD+o2 model. In [yang2014overlapping] it was found that, in real networks, community intersections can be significantly denser than the individual communities. In Figure 6 we confirm that the DBLP, Amazon, and YouTube networks all exhibit this phenomenon. Moreover, we compare these densities with those of the corresponding ABCD+o2 graphs. Aside from confirming that the model can produce behaviour similar to the real networks, we also find that the dimension of the reference layer affects the density of both the communities and the intersections. Moreover, we see that a low-dimensional reference layer yields a better approximation of the real data in the case of Amazon, whereas a high-dimensional reference layer yields a better approximation in the cases of DBLP and YouTube. In the case of DBLP, the difference in distribution between dimensions is partially attributable to the difference in the empirical $\rho$ (recall with Table 3 that the low dimension graphs could not achieve a large correlation. Note that the most similar dimension for each real graph is consistent with the experiment presented in Figure 3.
Figure 6: Distribution of the density for pairs of communities with non-zero overlap. The shaded region corresponds to the 25th to 75th quartiles, and the whiskers correspond to the maximum and minimum values. We restrict to overlaps of size at least $10$ due to the discrete and unpredictable nature of small communities.
@@
Finally, we analyze how strongly nodes are connected to their communities. There are many measures for community association strength, some of which we recently analyzed in [Betastar2025]. For the sake of simplicity, the measure we will use here is the internal edge fraction $\mathrm{IEF}(v,C)$, defined as
@@
Figure 7: Five largest internal edge fractions, sorted in descending order, grouped by the number of communities. A pattern is consistent between DBLP and ABCD+o2 , namely, that for nodes belonging to $\ell$ communities there are $\ell$ large, but not equal, internal edge fraction values.
@@ 4 Benchmarking Community Detection Algorithms
Leiden [leiden19]: a greedy algorithm that attempts to optimize the modularity function. Note that this algorithm returns a partition, and we use it merely as a baseline to compare with algorithms that attempt to find overlapping communities.
Edge Clustering [kim2015detecting]: an edge-partitioning algorithm that translates to overlapping clusters of nodes. Pairs of edges are measured based on similarity of neighbourhoods, and these similarity measures dictate the order in which edge-communities merge. As edge-communities merge, the modularity is tracked on the line-graph, and the maximum modularity attained yields the edge-communities, which in turn yields overlapping node-communities.
Ego-Split [epasto2017ego]: a method which finds overlapping clusters in a graph $G$ by applying a partitioning algorithm such as Leiden to an auxiliary graph $G^{\prime}$ and mapping the resulting partition onto $G$. The auxiliary graph $G^{\prime}$ is constructed from $G$ by creating multiple copies, or “egos”, of each node based on its neighbourhood.
Ego-Split+IEF [Betastar2025]: the same algorithm as Ego-Split, but with a post-processing step that re-assigns nodes to communities based on the IEF measure.
@@
In the first experiment, we compare the accuracy of each algorithm with respect to the ground-truth communities. The measure we use to determine accuracy is the overlapping Normalized Mutual Information (oNMI) measure [mcdaid2011normalized]. For this experiment we fix the number of nodes $n=5,000$, the minimum degree $\delta=5$, the maximum degree $\Delta=100$, the minimum community size $s=50$, and the maximum community size $S=500$. The parameters of ABCD+o2 with the most influence on the quality of detection algorithms are $\xi$ (the level of noise) and $\eta$ (the average number of communities a non-outlier is part of), and we consider every combination of $\xi\in\{0.1,0.2,\dots,0.6\}$ and $\eta\in\{1,1.25,\dots,2.5\}$. We fix the remaining parameters according to the YouTube graph, namely, $\gamma=1.87$, $\beta=2.13$, and $\rho=0.37$.
@@
Figure 10: The runtime of four clustering algorithms on ABCD+o2 graphs with 1,000 to 15,000 nodes.